
GDSR infection removal complete, evaluate DELS..
Source: http://2012portal.blogspot.com/2019/06/blog-post_5.html

GDSR infection removal complete, evaluate DELS..
Source: http://2012portal.blogspot.com/2019/06/blog-post_5.html

EXMOSS Pandora sequence approved..
Source: http://2012portal.blogspot.com/2019/05/blog-post.html

At the moment of the last Archon invasion in January 1996, the dark forces have created three main dark wormhole entry points to collapse the planetary Light grid…
Source: http://2012portal.blogspot.com/2019/05/situation-update.html

Pandora / GCOM in progress, ANNABELLA fully operational, internal YXR flow / multiphasic recovery in progress. Surface GDS coding and VTXPOS requirements fully enforced. Projects 310 and 061 in progress. Minimum M / HP requirements met, SD requirements not met…
Source: http://2012portal.blogspot.com/2019/05/state-of-mission-report-april-2019.html

http://bit.ly/2LjVVYe They have ordered Jesuits to put key Templar sacred sites on fire immediately, as Templars were the main force recovering fragments of Isis mysteries and bringing them into the awareness of the Western civilization…
Source: http://2012portal.blogspot.com/2019/04/the-high-priestess.html

Since ANNABELLA.exe, minimum hyperphase requirements and conversion streams are all finalized and stable, their status will not be reported in future updates unless necessary. VTXPOS requirements are more strictly enforced, limiting towards full INT YXR flow recovery…
Source: http://2012portal.blogspot.com/2019/04/project-501-update.html

VTXPOS minimum system requirements strictly enforced..
Source: http://2012portal.blogspot.com/2019/04/blog-post_76.html

Our Soul Families workshop in Budapest was extremely successful…
Source: http://2012portal.blogspot.com/2019/04/soul-families-workshop-report-and.html

LSIP BNF ANNABELLA.exe v1.3 is fully operational and stable. Interferences into INT YXR and GBN flow, PT intersection failure, recovery in progress. Project 310 minimum requirements met, Project 061 activated. SEP non-local overconversion! Minimum hyperphase requirements are finalized at PHI 1.5-3…
Source: http://2012portal.blogspot.com/2019/03/project-501-update.html

There are many things happening behind the scenes and if everything goes according to the plan, an update for the surface population will be posted on this blog in a few days…
Source: http://2012portal.blogspot.com/2019/03/a-short-notice-to-surface-population.html

Project 501 update expected 190331, SD grid recovery in progress..
Source: http://2012portal.blogspot.com/2019/03/blog-post_30.html

Do NOT contact trustproject@tutanota.com . That email address has been hijacked…
Source: http://2012portal.blogspot.com/2019/03/urgent-message-to-certain-group.html

Wonderland integrated into LSIP ANNABELLA…
Source: http://2012portal.blogspot.com/2019/01/wonderland-integrated-into-lsip.html

LSIP Wonderland protocols systemic surface failure, urgent evaluation, ASCOM briefing at 190107…
Source: http://2012portal.blogspot.com/2019/01/lsip-wonderland-protocols-systemic.html

Return of Light meditation videos have been created in more than 20 languages. Make them viral, spread them through your networks to assist in reaching the critical mass! English: https://www.youtube.com/watch?v=RW4ENEeT2uY&t=3s https://www.youtube.com/watch?v=a97NChtqiPs&feature=youtu…
Source: http://2012portal.blogspot.com/2019/01/return-of-light-meditation-update.html

Cobra in progress, Pandora in progress. Fatal security breach at 741, alpha timeline collapse. Beta recovery attempt failed, beta timeline collapse. GHVBN attempt, HP+/GBN-, EXC. Gamma timeline in progress. Several PB removal attempts with overall failure…
Source: http://2012portal.blogspot.com/2019/01/state-of-mission-report-2018.html

1. Use your own technique to bring you to a relaxed state of consciousness. 2. State your intent to use this meditation as a tool to bring the Light to the surface of the planet…
Source: http://2012portal.blogspot.com/2018/12/make-this-viral-return-of-light.html

PLOC has been confirmed, leading to DGRID exposure. LSIP is not yet supported due to unfavorable grid ratio. Minimum SDBN requirements are set at PHI 2.5-3.5, P>0.8, THETA>0.9, VTXPOS>2.0. Minimum hyperphase streaming requirements are set at beta (numex stream): PHI 1.5-3.5, P>0.9 THETA >0…
Source: http://2012portal.blogspot.com/2018/12/project-501-update.html